OpenStudy (anonymous):

How do I solve an equation by the square root property? Like: (2x-1)squared=53

OpenStudy (radar):

\[(2x-1)^{2}=53\] Perform the indicated squaring. \[4x ^{2}-4x+1=53\] Subtract 53 from both sides getting\[4x ^{2}-4x-52=0\] Divide thru by 4 getting\[x ^{2}-x-13=0\]

OpenStudy (radar):

Solve using quadratic formula.

OpenStudy (radar):

a=1, b=-1, and c=-13